OpenStudy (anonymous):

A cylindrical well is 20 meters deep and has a diameter of 1.5 meters. Approximately how many cubic meters of soil were dug out to make the well? (Use π = 3.14.) 11.78 cubic meters 35.33 cubic meters 121.20 cubic meters 141.30 cubic meters

OpenStudy (anonymous):

@TheSmartOne @pyxieyt

TheSmartOne (thesmartone):

\(\sf\Large V = \pi r^2 h\)

TheSmartOne (thesmartone):

What is 1.5/2 = ? Diameter/2 = radius and we need to find the radius :)

OpenStudy (anonymous):

0.75

TheSmartOne (thesmartone):

and h= 20 and \(\pi\) = 3.14 \(\sf\Large V = (3.14) \times (0.75)^2 \times 20 =?\)

OpenStudy (anonymous):

it is 1.76625

OpenStudy (anonymous):

i think

OpenStudy (anonymous):

@TheSmartOne

TheSmartOne (thesmartone):

nope, that isn't even an answer choice :/

TheSmartOne (thesmartone):

ok what is 0.75 x0.75 = ? you need to use a calculator

OpenStudy (anonymous):

okay 0.5625

OpenStudy (anonymous):

okay i got the answer now i read it wrong ts 35.33

OpenStudy (anonymous):

thank you

TheSmartOne (thesmartone):

that's correct :)

TheSmartOne (thesmartone):

anytime! :D